# Charges, complex structures, and perturbations of instantons

### v2

@@ Abstract.
Hermitian non-Kähler Einstein 4-manifolds have a quasi-locally conserved charge associated to spin-lowering via Killing spinors, and corresponding to a parameter of the moduli space. This charge is evaluated for all explicitly known examples. Infinitesimal Einstein deformations are shown to admit a closed 2-form that measures the perturbation to this charge. Provided certain boundary conditions hold, a curve of metrics in the moduli space passing through a Hermitian non-Kähler Einstein metric is shown to be conformally Kähler to second perturbative order. This result is applied to prove that infinitesimal Ricci-flat deformations of ALF Hermitian non-Kähler instantons are tangent to the moduli space of Hermitian instantons, and hence as a consequence that infinitesimal rigidity holds. The analog of this statement also holds for the case of compact Hermitian non-Kähler Einstein metrics.
@@ 1. Introduction
In Euclidean quantum gravity, the dominant contribution to the gravitational path integral is expected to come from metrics near the minima of the Euclidean action [23], that is from gravitational instantons and small deformations of them. In the one-loop expansion, deformations of a classical solution $(M,g_{ab})$ represent quantum corrections [25], and in the case of infinitesimal Einstein deformations, i.e. perturbations satisfying the linearized Einstein equations, they provide information about the moduli space $\mathscr{E}(M)$ of instantons, describing the structure of the tangent space of $\mathscr{E}(M)$ at $g_{ab}$ and allowing to analyze questions such as integrability, infinitesimal rigidity [4] and spectral stability (e.g. [18, 11]); cf. [9] for the compact case. This paper is concerned with uncovering some special properties of this class of deformations.
The study of the moduli space is further motivated by the recent discovery of the Chen-Teo instanton [16], a Ricci-flat, asymptotically flat (AF) solution that gives a counter-example to the Euclidean black hole uniqueness conjecture [22, 24] and contributes, in addition to flat space and to the Schwarzschild and Kerr instantons, to the possible semi-classical states of the gravitational field with AF boundary conditions [32]. In principle, further unknown gravitational instantons could exist, and the classification problem remains largely open, similarly to the compact case where all known examples have special holonomy [9]; see [3].
For the vacuum Einstein equations with or without cosmological constant, other boundary conditions of interest in quantum gravity, in addition to AF, include asymptotically locally flat (ALF), asymptotically locally Euclidean (ALE), and compact manifolds. All known instantons are (at least locally [^1] Our definition of instantons here allows a non-vanishing cosmological constant. This includes the 4-sphere $S^{4}$, which does not admit a global complex structure.) Hermitian [2]. Furthermore, they are conformally Kähler, which means [19] that there is a metric $\hat{g}_{ab}=\Omega^{2}g_{ab}$ and a self-dual 2-form $\hat{\omega}_{ab}$ which is parallel under the Levi-Civita connection ^$\hat{\nabla}_{a}$ of $\hat{g}_{ab}$,
^$\hat{\nabla}_{a}\hat{\omega}_{bc}=0.$ (1.1)
In this paper, we shall see that for Hermitian instantons, if the part of the Weyl tensor that has the same orientation as the complex structure, say $\Psi_{ABCD}$ in 2-spinor notation [39, 40] (cf. section 2.1 for conventions), is non-vanishing, then one of the moduli corresponds to a special “charge” associated to the gravitational field. This is because the Kähler form is actually the spin-lowered Weyl spinor
$\hat{\omega}_{ab}=\Psi_{ABCD}K^{CD}\epsilon_{A^{\prime}B^{\prime}},$ (1.2)
where $K^{AB}$ is a real solution to the 2-index twistor equation, see section 3.1 below. The gravitational charges referred to above are then the periods of the closed 2-form (1.2) over 2-cycles $S$ in $M$,
$Q[S]=\int_{S}\hat{\omega}.$ (1.3)
Since ${\rm d}\hat{\omega}=0$, (1.3) is invariant under changes of representative of the homology class $[S]$, so the charge is both geometrical (it depends on the curvature of $g_{ab}$) and topological (it depends on the homology groups of $M$).
The above result is the Riemannian analog of a Lorentzian conservation law first found by Jordan, Ehlers, and Sachs [29], and then reformulated in [1] in terms of spin-lowering [40, §6.4] for the Kerr black hole, where the charge (1.3) corresponds to the mass parameter and the conservation law is intuitively clear: 2-spheres surrounding the black hole are non-contractible, and the mass contained within any two of them is the same. However, for gravitational instantons, the evaluation of (1.3) and its interpretation are considerably more subtle, since the topology may be substantially different from the Kerr topology $\mathbb{R}^{2}\times S^{2}$, and it is not obvious how to find the 2-cycles (i.e. the 2-spheres in Kerr). As we shall see, examples of (1.3) include: the NUT-charge/mass of the ALF Taub-bolt instanton on $\mathbb{CP}^{2}\setminus\{\rm pt.\}$, the only free parameter of the ALE Eguchi-Hanson instanton on $T^{*}\mathbb{CP}^{1}$, the cosmological constant for the compact instantons $\mathbb{CP}^{2}$ and $\mathbb{CP}^{2}\#\overline{\mathbb{CP}}^{2}$, and, remarkably, the two parameters of the AF Chen-Teo instanton on $\mathbb{CP}^{2}\setminus S^{1}$. The fact that the two parameters of Chen-Teo are integrals of the Kähler form is to be contrasted with the Kerr instanton, where (1.3) gives only the mass. Intuitively, the underlying reason is that the Chen-Teo instanton is in fact a “double” object (cf. [17, §5]), and each of the two 2-cycles carries a gravitational charge (1.3).
We shall prove that infinitesimal Einstein deformations of Hermitian instantons, with or without cosmological constant $\lambda$ and with any asymptotic structure, admit a 2-form $\delta\hat{\omega}_{ab}$ (given by (3.21)) which is closed, ${\rm d}\,\delta\hat{\omega}=0$, and whose integral over 2-cycles measures the perturbation to (1.3),
$\delta{Q}[S]=\int_{S}\delta\hat{\omega}.$ (1.4)
Moreover, the result (1.4) is gauge invariant (i.e. invariant under infinitesimal diffeomorphisms [^2] Gauge-invariance is important since $\mathscr{E}(M)$ involves a quotient by ${\rm Diff}(M)$.; see eq. (3.13) below). While (1.4) is valid for any boundary conditions, a particular application concerns the ALF Ricci-flat but non-Kähler case (the only possibilities being Kerr, Taub-bolt, and Chen-Teo [10, 35]), where infinitesimal Einstein deformations are expected to be perturbations with respect to the moduli parameters [4]. The formula (1.4) then gives the perturbation to one of the moduli of the Kerr, Taub-bolt, and Chen-Teo instantons.
The above result for perturbations is inspired by the conservation law for the Lorentzian case found by Fackerell [20] (cf. also [21]), which was formulated in terms of spin-lowering in [1] and used to argue that the analog of (1.4) gives the linearized mass of a perturbed Kerr black hole [1, Theorem 5.1].
Finally, while the conservation of (1.4) is the Riemannian analog of the Lorentzian result mentioned before, in this paper we shall also prove the following stronger result. For the compact and ALF cases, if $g_{ab}(s)$ is a smooth curve of Einstein metrics in $\mathscr{E}(M)$, with $g_{ab}(0)$ Hermitian, we shall prove that
Here, we have a metric $\hat{g}_{ab}(s)=\Omega{(s)}^{2}g_{ab}(s)$ with Levi-Civita connection ^$\hat{\nabla}^{(s)}_{a}$, and $\hat{\omega}_{ab}(s)$ is a normalized eigenform of its self-dual Weyl tensor with eigenvalue $\Omega(s)$. In view of (1.1), the result (1.5) is gauge invariant. Note that, while the conservation law for (1.4) follows from (1.5) (by antisymmetrizing the indices), the result (1.5) is stronger.
The interpretation of (1.5) is as follows. The complex structure corresponding to $\hat{\omega}_{ab}(s)$ is parallel to second order, and hence the curve of metrics $g_{ab}(s)$ in $\mathscr{E}(M)$ is conformally Kähler and in particular Hermitian to second order, with charges $(Q+s\,\delta Q)[S]$. This is consistent with the recent result of Biquard, Gauduchon, and LeBrun [12], showing that local rigidity holds for Hermitian, toric, ALF instantons, i.e. that Ricci-flat metrics in a suitable neighbourhood of a Hermitian, toric, ALF Ricci-flat instanton are Hermitian. Local rigidity, however, does not imply infinitesimal rigidity, in the sense that an ALF Ricci-flat perturbation is tangent to the moduli space of Hermitian instantons. As an application of (1.5), we are now able to prove that infinitesimal rigidity holds.
Recall that an Einstein metric is said to be integrable if any linearized Einstein perturbation is tangent to a curve of Einstein metrics, see [9]. It is an open problem whether integrability holds in general for Einstein metrics. It was conjectured [4, Conjecture 1.7] that ALF Hermitian non-Kähler instantons are integrable in the above sense. From (1.5), we have that any infinitesimal Einstein deformation is Hermitian, and since the moduli space is smooth and finite-dimensional, we have the following theorem.
@@
Overview
In section 2 we introduce some background material and set our notation and conventions. In section 3 we prove an identity for generic Riemannian 4-manifolds (Prop. 3.1), and use it to derive the closed 2-form $\delta\hat{\omega}_{ab}$ and the linearized charge (1.4); see Theorem 3.3. Section 4 contains many examples of the charge (1.3) for different gravitational instantons (AF, ALF, ALE, compact). In section 5, we prove (1.5) using a divergence identity (Lemma 5.1) based on LeBrun’s work [34], see Theorem 5.3, and in section 5.3 we use this to prove Theorem 1.1.
@@ 2.1. Notation and conventions
Throughout the article, $(M,g_{ab})$ denotes a 4-dimensional, orientable Riemannian manifold, with metric signature (++++)(++++)( + + + + ), Levi-Civita connection $\nabla_{a}$, volume form $\epsilon_{abcd}$, and Hodge star operator *. We shall use, locally, the 2-spinor formalism [39, 40], adapted to Riemannian signature. We do not assume a global spin structure.
The (locally defined) unprimed and primed spinor bundles are $\mathbb{S}$ and $\mathbb{S}^{\prime}$. The symplectic forms $\epsilon_{AB}$ and $\epsilon_{A^{\prime}B^{\prime}}$ and their inverses are used to raise and lower spinor indices. An unprimed spin dyad is $(o^{A},\iota^{A})$ with $\epsilon_{AB}o^{A}\iota^{B}=1$ and $\iota^{A}=o^{\dagger A}$, and a primed spin dyad is $(\alpha^{A^{\prime}},\beta^{A^{\prime}})$ with $\epsilon_{A^{\prime}B^{\prime}}\alpha^{A^{\prime}}\beta^{B^{\prime}}=1$ and $\beta^{A^{\prime}}=\alpha^{\dagger A^{\prime}}$. Defining the null tetrad $\ell^{a}=o^{A}\alpha^{A^{\prime}}$, $n^{a}=\iota^{A}\beta^{A^{\prime}}$, $m^{a}=o^{A}\beta^{A^{\prime}}$, $\tilde{m}^{a}=\iota^{A}\alpha^{A^{\prime}}$ (see e.g. [39, (3.1.14)]), we have $n^{a}=\bar{\ell}^{a}$ and $\tilde{m}^{a}=-\bar{m}^{a}$, and the non-trivial inner products are $\ell^{a}n_{a}=1=-m^{a}\tilde{m}_{a}$. The real and imaginary parts of the null tetrad define four real vectors $\{e^{a}_{\bf a}\}$, ${\bf a}=0,...,3$, as follows: $e^{a}_{0}=\frac{1}{\sqrt{2}}(\ell^{a}+n^{a})$, $e^{a}_{1}=i\frac{1}{\sqrt{2}}(\ell^{a}-n^{a})$, $e^{a}_{2}=\frac{1}{\sqrt{2}}(m^{a}-\tilde{m}^{a})$, $e^{a}_{3}=i\frac{1}{\sqrt{2}}(m^{a}+\tilde{m}^{a})$. (This is the analog of [39, Eq. (3.1.20)].) The dual basis, denoted $e_{a}^{\bf a}$, is defined by $e_{a}^{\bf a}e^{a}_{\bf b}=\delta^{\bf a}_{\bf b}$. A computation gives:
@@
Thus the reality conditions on the null tetrad imposed above, which are in turn defined by the Riemannian spinor conjugation $\dagger$, give the correct signature (++++)(++++)( + + + + ) for the metric. (Note that an analogous calculation but with Lorentzian reality conditions gives signature (+)(+---)( + - - - ), cf. footnote in [39, p. 121].)
Let $\Lambda^{2}$ be the space of 2-forms in $M$. This splits under * as $\Lambda^{2}=\Lambda^{2}_{+}\oplus\Lambda^{2}_{-}$, where $\Lambda^{2}_{\pm}$ corresponds to the eigenvalue ±$\pm 1$ of *. Elements of $\Lambda^{2}_{+}$ are self-dual 2-forms, and elements of $\Lambda^{2}_{-}$ are anti-self-dual 2-forms. Our convention for the volume form is $\epsilon_{0123}=+1$, i.e. $\epsilon_{abcd}=4!e^{0}_{[a}e^{1}_{b}e^{2}_{c}e^{3}_{d]}$, or equivalently $\boldsymbol{\epsilon}=+e^{0}\wedge e^{1}\wedge e^{2}\wedge e^{3}$. With this convention, one can check that a basis of real self-dual 2-forms is:
@@
The 2-forms (2.1) give an orthonormal basis of $\Lambda^{2}_{+}$. In terms of spinors, this can be phrased in terms of the real ‘metric’ on $\mathbb{S}^{2}=\mathbb{S}\odot\mathbb{S}$ given by
@@
cf. [40, Eq. (8.3.2)], with inverse $I^{ABCD}$. Writing (2.1) as $Z^{i}_{ab}=\zeta^{i}_{AB}\epsilon_{A^{\prime}B^{\prime}}$ ($i=1,2,3$), the spinors $\zeta^{i}_{AB}$ are real, symmetric, and satisfy
@@
$A_{a}={}$ $\Gamma_{a}{}^{3}{}_{1}=-(o^{B}\nabla_{a}o_{B}+\iota^{B}\nabla_{a}\iota_{B}),$ (2.5a)
$B_{a}={}$ $\Gamma_{a}{}^{3}{}_{2}=i(o^{B}\nabla_{a}o_{B}-\iota^{B}\nabla_{a}\iota_{B}),$ (2.5b)
$C_{a}={}$ $\Gamma_{a}{}^{2}{}_{1}=-2i\,\iota^{B}\nabla_{a}o_{B}.$ (2.5c)
@@
We stress that this is the opposite of the convention used by Penrose and Rindler [39, 40] (compare to [39, Eq. (4.2.31)]). This means that the Riemann tensor here is minus the Riemann tensor in [39, 40], and similarly the curvature spinors here are minus the curvature spinors in [39, 40]. This convention has been adopted to match the sign convention of the determinant of the Weyl tensor in references [42, 34, 12], which we shall invoke in section 5.
@@
where $\Box_{AB}=\nabla_{A^{\prime}(A}\nabla^{A^{\prime}}_{B)}$ and $\Box_{A^{\prime}B^{\prime}}=\nabla_{A(A^{\prime}}\nabla^{A}_{B^{\prime})}$ (compare to [39, Eq. (4.9.15)]). In (2.7), $R=g^{ac}R_{abc}{}^{b}$ is the scalar curvature, and $\Psi_{ABCD}$ and $\tilde{\Psi}_{A^{\prime}B^{\prime}C^{\prime}D^{\prime}}$ are the Weyl curvature spinors. The Weyl tensor is $W_{abcd}=W^{+}_{abcd}+W^{-}_{abcd}$, where

### v3

@@ Abstract.
The main result of this paper is that infinitesimal rigidity of the moduli space for Hermitian gravitational instantons holds. An important step in the proof is to show that, provided certain boundary conditions hold, a curve of Riemannian metrics passing through a Hermitian non-Kähler Einstein metric is conformally Kähler to second perturbative order. This uses ideas of Wu and LeBrun. Hermitian non-Kähler Einstein 4-manifolds have a quasi-locally conserved charge, which is shown to correspond to a parameter of the moduli space. This charge is evaluated for all explicitly known examples. It follows from the proof of our main theorem that infinitesimal Einstein deformations admit a closed 2-form that measures the perturbation in the moduli parameter.
@@ 1. Introduction
A gravitational instanton is a compact, or complete and asymptotically flat, Einstein or Ricci-flat 4-manifold. The concept of a gravitational instanton was introduced by Hawking [37] as an analog of instantons in quantum field theory [53]. While the notion of gravitational instanton is sometimes taken to imply special geometry such as the half-flat condition we do not, unless otherwise stated, impose such restrictions here. With suitable fall-off conditions on curvature, the asymptotically flat instantons can be classified depending on volume growth at infinity, the ones relevant here are asymptotically locally Euclidean (ALE, with quartic volume growth) and asymptotically flat (ALF, with cubic volume growth), but there are examples with quadratic, linear and even non-polynomical volume growth, cf. [52] and references therein. ALF instantons are asymptotic to a circle bundle over $S^{2}$ at infinity. ALF instantons where this circle bundle is trivial are called asymptotically flat (AF).
Important examples of gravitational instantons include the Euclidean signature versions of the Schwarzschild and Kerr black holes, the half-flat Taub-NUT and Eguchi-Hansen spaces, Chen-Teo and the compact Page and Chen-LeBrun-Weber spaces. See [4, 43] and references therein. Methods for construction of new instantons include resolution of singularities, gluing, the Gibbons-Hawking method, and inverse scattering. The Gibbons-Hawking method allows one to construct large families of half-flat ALF or ALE instantons, which are in general $S^{1}$-symmetric. The inverse scattering, or Belinskii-Zakharov method, was used to construct the Chen-Teo instanton [20], an ALF instanton which is Hermitian with respect to only one orientation [2]. Symmetry reduction of the Einstein equation leads to a harmonic map equation and this has been exploited in the work of Li and Sun [46], following the approach of Kunduri and Lucietti [41], to construct infinite families of toric AF instantons.
As proved by Derdzinski [24], Hermitian instantons are conformal to extremal Kähler spaces. This was used by LeBrun to provide a complete classification of compact Hermitian-Einstein instantons [43]. Biquard and Gauduchon obtained a complete classification of toric, Hermitian non-Kähler ALF instantons [14]. Recently the classification was extended by Li [45] to Hermitian ALF and ALE instantons. The classification of Hermitian ALE instantons is not complete, but requires some additional conditions on the structure at infinity. In the toric case, recent work by Lucietti and one of the authors provides a classification of Hermitian toric ALE instantons without additional conditions at infinity [10]. By the classification results for Hermitian instantons, the new examples of Li and Sun [46] are necessarily algebraically general. In [4], one of the authors and collaborators provided steps towards a classification of $S^{1}$-symmetric instantons. In particular, this result shows that an $S^{1}$-symmetric instanton with the same topology as the Kerr, Chen-Teo, or Taub-bolt spaces necessarily in the known moduli space. In view of the results of Li and Sun, the general classification problem is wide open, even for the case of toric instantons.
The work mentioned above gives many examples of moduli spaces of gravitational instantons. The known moduli spaces of ALF instantons on a manifold with given topology have dimension at most 2111Our definition of moduli space does not factor out the scaling, in contrast to e.g. [44]. As we shall see below, for the Hermitian instantons, the moduli parameters correspond to conserved charges.
Local rigidity for Hermitian ALF instantons has been recently proved by Biquard, Gauduchon and LeBrun [16]. This result is analogous the local rigidity theorems valid for Calabi-Yau spaces [23]. However, it is a priori possible that the moduli space of gravitational instantons fails to be integrable [12] in the sense that there could exist infinitesimal Einstein deformations that are not tangent to curves in the moduli space, cf. also discussion in section 5. Thus the result on local rigidity leaves open the question of infinitesimal rigidity and integrability of Hermitian instantons. In [5], we conjectured that integrability holds for Hermitian ALF instantons. The following theorem, which is the main result of this paper, settles this question.
@@
The proof makes use of the fact that a Hermitian Einstein 4-manifold is conformally Kähler. Let $\hat{g}_{ab}=\Omega^{2}g_{ab}$ be the Kähler metric conformal to $g_{ab}$, with Kähler form $\hat{\omega}_{ab}$, parallel with respect to the Levi-Civita connection $\hat{\nabla}_{a}$ of $\hat{g}_{ab}$,
$\hat{\nabla}_{a}\hat{\omega}_{bc}=0.$ (1.1)
The Kähler form is actually the spin-lowered Weyl spinor
$\hat{\omega}_{ab}=\Psi_{ABCD}K^{CD}\epsilon_{A^{\prime}B^{\prime}},$ (1.2)
where $K^{AB}$ is a real solution to the 2-index twistor equation, see section 3.1 below. In fact, we prove, using an argument inspired by the works of Wu [55] and LeBrun [44], that for the compact and ALF cases, the linearized version of (1.1) is also true.
Lemma 1.2.
Let $g_{ab}(s)$ be a smooth curve of Riemannian metrics on $M$, where $g_{ab}=g_{ab}(0)$ is Hermitian-Einstein and $\frac{{\rm d}}{{\rm d}s}g_{ab}(s)|_{s=0}$ is an infinitesimal Einstein deformation. Let $\hat{g}_{ab}(s)=\Omega{(s)}^{2}g_{ab}(s)$, with Levi-Civita connection $\hat{\nabla}^{(s)}_{a}$, where $\hat{\omega}_{ab}(s)$ is a normalized eigenform of the self-dual Weyl tensor corresponding to the positive eigenvalue $\Omega(s)$ [^2] Since $g_{ab}(0)$ is Hermitian, there is always one positive eigenvalue, and by continuity this holds for $s$ close to zero.. Then
Remark 1.3.
Note there are no assumptions on the higher order terms of $g_{ab}(s)$ in Lemma 1.2.
The periods of $\hat{\omega}_{ab}$ over 2-cycles $S$ in $M$ defines a gravitational charge
$Q[S]=\int_{S}\hat{\omega}$. (1.4)
This fact is the Riemannian analog of a Lorentzian conservation law first found by Jordan, Ehlers, and Sachs [38], and then reformulated in [1] in terms of spin-lowering [51, §6.4] for the Kerr black hole, where the charge (1.4) corresponds to the mass parameter. The resulting conservation law has an intuitively clear meaning, namely that the 2-spheres surrounding the black hole are non-contractible, and the mass contained within any two of them is the same.
However, for gravitational instantons, the evaluation of (1.4) and its interpretation are considerably more subtle, since the topology may be substantially different from the Kerr topology $\mathbb{R}^{2}\times S^{2}$, and it is not obvious how to find the 2-cycles (i.e. the 2-spheres in Kerr). As we shall see, examples of (1.4) include the NUT-charge/mass of the ALF Taub-bolt instanton on $\mathbb{CP}^{2}\setminus\{\rm pt.\}$, the only free parameter of the ALE Eguchi-Hanson instanton on $T^{*}\mathbb{CP}^{1}$, the cosmological constant for the compact instantons $\mathbb{CP}^{2}$ and $\mathbb{CP}^{2}\#\overline{\mathbb{CP}}{}^{2}$, and, remarkably, the two parameters of the AF Chen-Teo instanton on $\mathbb{CP}^{2}\setminus S^{1}$. The fact that the two parameters of Chen-Teo are integrals of the Kähler form is to be contrasted with the Kerr instanton, where (1.4) gives only the mass. Intuitively, the underlying reason is that the Chen-Teo instanton is in fact a ``double'' object (cf. [21, §5]), and each of the two 2-cycles carries a gravitational charge (1.4).
It follows from Lemma 1.2 that the linearized version of (1.4) also yields a conserved charge,
$\delta{Q}[S]=\int_{S}\delta\hat{\omega}$. (1.5)
Moreover, (1.5) is gauge invariant (i.e. invariant under infinitesimal diffeomorphisms; see eq. (3.13) below). While (1.5) is valid for any boundary conditions, a particular application concerns the ALF Ricci-flat but non-Kähler case (the only possibilities being Kerr, Taub-bolt, and Chen-Teo [14, 45]). By Theorem 1.1, infinitesimal Einstein deformations of a Hermitian ALF instanton are perturbations with respect to the moduli parameters, and hence the formula (1.5) gives the perturbation to one of the moduli of the Kerr, Taub-bolt, and Chen-Teo instantons.
The above result for perturbations is inspired by the conservation law for the Lorentzian case found by Fackerell [25], cf. also [26, 1]. The analog of (1.5) for the Lorentzian case gives the linearized mass of a perturbed Kerr black hole [1, Theorem 5.1].
1.1. Overview
In section 2 we introduce some background material and set our notation and conventions. In section 3 we prove an identity for generic Riemannian 4-manifolds (Prop. 3.1), and use it to derive the closed 2-form $\delta\hat{\omega}_{ab}$ and the linearized charge (1.5); see Theorem 3.3. Section 4 contains many examples of the charge (1.4) for different gravitational instantons (AF, ALF, ALE, compact). In section 5, we prove (1.3) using a divergence identity (Lemma 5.1) based on LeBrun's work [44], see Theorem 5.3, and in section 5.3 we use this to prove Theorem 1.1.
@@ 2.1. Notation and conventions
Throughout the article, $(M,g_{ab})$ denotes a 4-dimensional, orientable Riemannian manifold, with metric signature $(++++)$, Levi-Civita connection $\nabla_{a}$, volume form $\epsilon_{abcd}$, and Hodge star operator *. We shall use, locally, the 2-spinor formalism [50, 51], adapted to Riemannian signature. We do not assume a global spin structure.
The (locally defined) unprimed and primed spinor bundles are $\mathbb{S}$ and $\mathbb{S}^{\prime}$. The symplectic forms $\epsilon_{AB}$ and $\epsilon_{A^{\prime}B^{\prime}}$ and their inverses are used to raise and lower spinor indices. An unprimed spin dyad is $(o^{A},\iota^{A})$ with $\epsilon_{AB}o^{A}\iota^{B}=1$ and $\iota^{A}=o^{\dagger A}$, and a primed spin dyad is $(\alpha^{A^{\prime}},\beta^{A^{\prime}})$ with $\epsilon_{A^{\prime}B^{\prime}}\alpha^{A^{\prime}}\beta^{B^{\prime}}=1$ and $\beta^{A^{\prime}}=\alpha^{\dagger A^{\prime}}$. Defining the null tetrad $\ell^{a}=o^{A}\alpha^{A^{\prime}}$, $n^{a}=\iota^{A}\beta^{A^{\prime}}$, $m^{a}=o^{A}\beta^{A^{\prime}}$, $\tilde{m}^{a}=\iota^{A}\alpha^{A^{\prime}}$ (see e.g. [50, (3.1.14)]), we have $n^{a}=\bar{\ell}^{a}$ and $\tilde{m}^{a}=-\bar{m}^{a}$, and the non-trivial inner products are $\ell^{a}n_{a}=1=-m^{a}\tilde{m}_{a}$. The real and imaginary parts of the null tetrad define four real vectors $\{e^{a}_{\bf a}\}$, ${\bf a}=0,...,3$, as follows: $e^{a}_{0}=\frac{1}{\sqrt{2}}(\ell^{a}+n^{a})$, $e^{a}_{1}=i\frac{1}{\sqrt{2}}(\ell^{a}-n^{a})$, $e^{a}_{2}=\frac{1}{\sqrt{2}}(m^{a}-\tilde{m}^{a})$, $e^{a}_{3}=i\frac{1}{\sqrt{2}}(m^{a}+\tilde{m}^{a})$. (This is the analog of [50, Eq. (3.1.20)].) The dual basis, denoted $e_{a}^{\bf a}$, is defined by $e_{a}^{\bf a}e^{a}_{\bf b}=\delta^{\bf a}_{\bf b}$. A computation gives:
@@
Thus the reality conditions on the null tetrad imposed above, which are in turn defined by the Riemannian spinor conjugation $\dagger$, give the correct signature $(++++)$ for the metric. (Note that an analogous calculation but with Lorentzian reality conditions gives signature (+)(+---), cf. footnote in [50, p. 121].)
Let $\Lambda^{2}$ be the space of 2-forms in $M$. This splits under * as $\Lambda^{2}=\Lambda^{2}_{+}\oplus\Lambda^{2}_{-}$, where $\Lambda^{2}_{\pm}$ corresponds to the eigenvalue $\pm 1$ of *. Elements of $\Lambda^{2}_{+}$ are self-dual 2-forms, and elements of $\Lambda^{2}_{-}$ are anti-self-dual 2-forms. Our convention for the volume form is $\epsilon_{0123}=+1$, i.e. $\epsilon_{abcd}=4!e^{0}_{[a}e^{1}_{b}e^{2}_{c}e^{3}_{d]}$, or equivalently $\boldsymbol{\epsilon}=+e^{0}\wedge e^{1}\wedge e^{2}\wedge e^{3}$. With this convention, one can check that a basis of real self-dual 2-forms is:
@@
The 2-forms (2.1) give an orthonormal basis of $\Lambda^{2}_{+}$. In terms of spinors, this can be phrased in terms of the real `metric' on $\mathbb{S}^{2}=\mathbb{S}\odot\mathbb{S}$ given by
@@
cf. [51, Eq. (8.3.2)], with inverse $I^{ABCD}$. Writing (2.1) as $Z^{i}_{ab}=\zeta^{i}_{AB}\epsilon_{A^{\prime}B^{\prime}}$ ($i=1,2,3$), the spinors $\zeta^{i}_{AB}$ are real, symmetric, and satisfy
@@
$\Gamma_{a}{}^{3}{}_{1}$={} $-(o^{B}\nabla_{a}o_{B}+\iota^{B}\nabla_{a}\iota_{B}),$ (2.5a)
$\Gamma_{a}{}^{3}{}_{2}$={} $i(o^{B}\nabla_{a}o_{B}-\iota^{B}\nabla_{a}\iota_{B}),$ (2.5b)
$\Gamma_{a}{}^{2}{}_{1}$={} $-2i\,\iota^{B}\nabla_{a}o_{B}.$ (2.5c)
@@
We stress that this is the opposite of the convention used by Penrose and Rindler [50, 51] (compare to [50, Eq. (4.2.31)]). This means that the Riemann tensor here is minus the Riemann tensor in [50, 51], and similarly the curvature spinors here are minus the curvature spinors in [50, 51]. This convention has been adopted to match the sign convention of the determinant of the Weyl tensor in references [55, 44, 16], which we shall invoke in section 5.
@@
where $\Box_{AB}=\nabla_{A^{\prime}(A}\nabla^{A^{\prime}}_{B)}$ and $\Box_{A^{\prime}B^{\prime}}=\nabla_{A(A^{\prime}}\nabla^{A}_{B^{\prime})}$ (compare to [50, Eq. (4.9.15)]). In (2.7), $R=g^{ac}R_{abc}{}^{b}$ is the scalar curvature, and $\Psi_{ABCD}$ and $\tilde{\Psi}_{A^{\prime}B^{\prime}C^{\prime}D^{\prime}}$ are the Weyl curvature spinors. The Weyl tensor is $W_{abcd}=W^{+}_{abcd}+W^{-}_{abcd}$, where